# The fate of poor man’s Majoranas in the long Kitaev chain limit

### v1

@@ Abstract
A minimal Kitaev chain, consisting of two quantum dots connected via a superconductor, can host highly localized near-zero-energy states, known as poor man’s Majoranas (PMMs). These states have been proposed as promising candidates to study Majorana bound states (MBSs) in a highly tunable setup. However, it is unclear whether and how PMMs observed in real systems are actually connected to the topological phase of the full Kitaev chain. Here, we study PMMs using a microscopic model and show that, in the long chain limit, not all PMMs are related to topological states. Rather, in long chains, some PMMs evolve into trivial highly localized low-energy states. We provide an analytical explanation for the occurrence of these states and show that there is no clear conductance signature that is able to distinguish PMMs that evolve into true topological states from PMMs that evolve into trivial states.
@@
Majorana bound states (MBSs) are quasiparticles that emerge at the boundaries or in vortices of topological superconductors Kitaev (2001) and, due to their non-Abelian exchange statistics Ivanov (2001), have been proposed as candidates to store and manipulate quantum information in a fault-tolerant way Kitaev (2003); Nayak et al. (2008); Elliott and Franz (2015). They were prominently described using the Kitaev chain Kitaev (2001), a simple spinless minimal model relying on $p$-wave superconductivity. Nanowires with strong spin-orbit interaction (SOI) proximitized by superconductors have in recent years been prominent systems to realize an effective Kitaev chain Lutchyn et al. (2010); Oreg et al. (2010); Stanescu et al. (2011); Mourik et al. (2012); Das et al. (2012); Deng et al. (2012); Laubscher and Klinovaja (2021). However, disorder has hindered the conclusive observation of MBSs in these heterostructures since disorder can result in signatures that mimic MBSs Kells et al. (2012); Lee et al. (2012); Rainis et al. (2013); Roy et al. (2013); Ptok et al. (2017); Liu et al. (2017); Moore et al. (2018a, b); Reeg et al. (2018); Vuik et al. (2019); Stanescu and Tewari (2019); Woods et al. (2019); Chen et al. (2019); Awoga et al. (2019); Prada et al. (2020); Yu et al. (2021); Das Sarma and Pan (2021); Valentini et al. (2021); Hess et al. (2021, 2023a, 2022). Due to the issues with disorder in nanowires, realizing an effective Kitaev chain as an array of quantum dots (QDs) has been proposed as an alternative platform to host MBSs with a high degree of tunability Sau and Sarma (2012); Leijnse and Flensberg (2012); Fulga et al. (2013); Sothmann et al. (2013); Li et al. (2014); Su et al. (2017); Tsintzis et al. (2022); Liu et al. (2022); Dvir et al. (2023); Sanches et al. (2023); Liu et al. (2023); Koch et al. (2023); Souto et al. (2023); Zatelli et al. (2023); Samuelson et al. (2024); Tsintzis et al. (2024); Pino et al. (2024a); Luna et al. (2024); Bordin et al. (2024a); Boross and Pályi (2024); Liu et al. (2024a); Vilkelis et al. (2024); Bozkurt et al. (2024); van Driel et al. (2024); Pino et al. (2024b); Geier et al. (2024); Ten Haaf et al. (2024); Kocsis et al. (2024); Cayao (2024); Miles et al. (2024); Cayao and Aguado (2024); Liu et al. (2024b); Alvarado et al. (2024); Svensson and Leijnse (2024); Álvaro Gómez-León et al. (2024); Pan et al. (2024); Pandey and Dagotto (2024); Liu et al. (2024c); Benestad et al. (2024); Luethi et al. (2024). In these proposals, the QDs are separated by superconducting sections, which couple the QDs through elastic cotunneling (ECT) and crossed Andreev reflection (CAR) Recher et al. (2001); Lesovik et al. (2001); Falci et al. (2001); Bouchiat et al. (2002); Feinberg (2003); Hofstetter et al. (2009); Herrmann et al. (2010). In such setups, an inhomogenous magnetic field or SOI can enable effective $p$-wave superconductivity using a conventional $s$-wave superconductor Sau and Sarma (2012); Leijnse and Flensberg (2012).
@@
It has been shown that a fully spin-polarized model consisting of only two QDs with a superconductor transmitting ECT and CAR between the QDs is sufficient to obtain states similar to MBSs Leijnse and Flensberg (2012). These states share most of their properties of the MBSs found in long Kitaev chains, i.e., they are at zero energy, separated from the excited states by a finite gap, and have non-Abelian exchange statistics. However, they exist only at finely tuned points of parameter space, called “sweet spots”, and therefore they lack topological protection. Due to this lack of protection MBSs in minimal Kitaev chains are often called “poor man’s Majoranas” (PMMs).
@@
Recent demonstrations of control over the strength of ECT and CAR Liu et al. (2022); Wang et al. (2022); Kürtössy et al. (2022); de Jong et al. (2023); Bordin et al. (2023); Wang et al. (2023); Bordin et al. (2024b) have spurred great interest in PMMs. Although originally derived using the spinless Kitaev chain, PMMs have also been studied in spinful models Fulga et al. (2013); Tsintzis et al. (2022); Liu et al. (2022); Sanches et al. (2023); Liu et al. (2024c); Souto et al. (2023); Zatelli et al. (2023); Samuelson et al. (2024); Tsintzis et al. (2024); Luna et al. (2024); Liu et al. (2024a); Vilkelis et al. (2024); Bozkurt et al. (2024); Geier et al. (2024); Ten Haaf et al. (2024); Kocsis et al. (2024); Miles et al. (2024); Liu et al. (2024b); Alvarado et al. (2024); Svensson and Leijnse (2024); Benestad et al. (2024); Luethi et al. (2024). It was shown in Ref. 73 that sweet spots with perfectly localized zero-energy PMMs can exist in an idealized spinful model, however, reaching these sweet spots requires complete control over many degrees of freedom. In more realistic models and experiments such control is not possible and so only “imperfect PMMs” are possible, i.e., states that are close to zero energy and highly – but not perfectly – localized.
In recent experiments, conductance signatures associated with PMMs have been found Dvir et al. (2023); Ten Haaf et al. (2024); Zatelli et al. (2023); Bordin et al. (2024a); van Driel et al. (2024). Whilst theoretical proposals for braiding PMMs also exist Boross and Pályi (2024); Tsintzis et al. (2024); Geier et al. (2024), these have not yet been demonstrated. Given that only imperfect PMMs are possible in realistic setups and experiments, it is important to clarify whether these states are actually connected to topological states in the Kitaev chain, or if they are simply well-localized near-zero-energy states that evolve into trivial states in the long chain limit.
@@
Figure 1: Microscopic model of a minimal Kitaev chain and the long chain limit. (a): Minimal Kitaev chain, consisting of three QDs. The left and right QDs, representing two Kitaev sites, have a chemical potential $\mu$, Zeeman energy $\Delta_{Z}$, and on-site Coulomb interaction $U$. The middle QD M has a superconducting pairing potential $\Delta$ and, due to the screening by the superconductor, we assume no Zeeman energy or Coulomb interaction. This middle QD models an Andreev bound state (ABS) that transmits the effective couplings between the two outer QDs. The hopping amplitude is $t$ and $U_{\mathrm{SOI}}(\Phi_{\mathrm{SOI}}/2)$ characterizes the SOI. (b): Long chain limit. Normal QDs are white, they are equivalent to the QDs L and R in (a) and each represents one normal site. The QDs with a pairing potential are blue and they are equivalent to the QD M in (a).
@@
In this Letter, we demonstrate that imperfect PMMs in a microscopic model of the minimal Kitaev chain are not necessarily connected to topological MBSs in the long chain limit. We explain analytically the origin of these imperfect PMMs and demonstrate that there is no clear conductance signature in a minimal Kitaev chain that can distinguish PMMs that evolve into true MBSs from PMMs that evolve into highly localized trivial states.
@@
Model: To model a minimal Kitaev chain microscopically we consider a system of three QDs, see Fig. 1(a), based on the model introduced in Ref. 40. The left and right QDs, modelling the two sites of a minimal Kitaev chain, have chemical potentials $\mu_{L}$ and $\mu_{R}$, respectively (if not stated otherwise, we will assume $\mu_{L}=\mu_{R}\equiv\mu$), a Zeeman energy $\Delta_{Z}$, and on-site Coulomb repulsion $U$. The middle QD is proximitized by a superconductor and models an Andreev bound state that effectively transmits ECT and CAR between the outer QDs Tsintzis et al. (2022); Liu et al. (2022); Miles et al. (2024); Souto et al. (2023); Liu et al. (2024c); Zatelli et al. (2023); Luna et al. (2024); Valentini et al. (2024). The middle QD also causes local Andreev reflection (LAR) and renormalizes other parameters of the outer QDs Luethi et al. (2024). This middle dot has a superconducting pairing potential $\Delta$ and a chemical potential $\mu_{M}$. Due to screening by the superconductor we will assume that the middle QD has no Zeeman energy and no Coulomb repulsion. We note, however, that it has been shown in Ref. 40 that, even without this assumption for the middle QD, PMMs persist in the system studied. The hopping between the QDs is determined by the hopping amplitude $t$ and the SOI angle $\Phi_{\mathrm{SOI}}$. The Hamiltonian describing the system is given by
@@
where $n_{j,\sigma}=d_{j,\sigma}^{\dagger}d_{j,\sigma}$, $d_{j,\sigma}^{\dagger}$ ($c_{\sigma}^{\dagger}$) creates a particle on dot $j\in\{L,R\}$ (the middle QD M), $\mu_{M}$ is the chemical potential on the middle QD, and the notation $\sigma\Delta_{Z}$ means $+\Delta_{Z}$ for $\sigma=\uparrow$ and $-\Delta_{Z}$ for $\sigma=\downarrow$. The chemical potentials are measured with respect to the chemical potential of the superconductor. Assuming that SOI is perpendicular to the magnetic field, the SOI matrix is given by
@@
Figure 2: Quality measures for true PMMs versus false PMMs. Energy difference $\Delta E$ [(a) and (d)], charge difference $\Delta Q_{j}$ [(b) and (e)], and the Majorana polarisation $M_{j}$ [(c) and (f)] for two parameter sets. The area circled in black is the ROT. Both sets of parameters lead to qualitatively equal results, and both lead to a ROT with highly localized near-zero-energy states, i.e., PMMs. However, only (a)-(c) are connected to true MBSs in the long chain limit, whereas (d)-(f) are connected to trivial states in the long chain limit, see Fig. 3. The parameters for (a)-(c) are $t/\Delta=0.42$, $\Phi_{\mathrm{SOI}}=0.26\pi$, $U=0$, and $\Delta_{Z}/\Delta=0.8$. The largest excitation gap in the ROT is $E_{\mathrm{ex}}/\Delta=0.103$. The parameters for (d)-(f) are $t/\Delta=0.99$, $\Phi_{\mathrm{SOI}}=0.44\pi$, $U=0$, and $\Delta_{Z}/\Delta=0.8$. The largest excitation gap in the ROT is $E_{\mathrm{ex}}/\Delta=0.084$. The ROT threshold values for all panels are $\Delta E_{\mathrm{th}}/\Delta=10^{-4}$, $\Delta Q_{\mathrm{th}}=0.05$, and $M_{\mathrm{th}}=0.05$.
@@
Region of threshold: A sweet spot is a point in parameter space that fulfills $\Delta E=0$, $\Delta Q_{L}=\Delta Q_{R}=0$, $M_{L}=M_{R}=1$, and $E_{\mathrm{ex}}>0$. The corresponding “perfect PMMs” to these sweet spots are completely localized and at zero energy. It was shown in Ref. Luethi et al. (2024) that, for the microscopic Hamiltonian of a minimal Kitaev chain in Eq. (1), no perfect sweet spots could be found. Instead, the notion of a “region of threshold” (ROT) was introduced. It is a region of parameter space where the PMM conditions are approximately realized, i.e.,
@@
where $\Delta E_{\mathrm{th}}$, $E_{\mathrm{ex,th}}$, $\Delta Q_{\mathrm{th}}$, and $M_{\mathrm{th}}$ are threshold values which need to be chosen. The corresponding “imperfect PMMs” have energies near zero and are highly – but not perfectly – localized. This highlights, however, that it is necessary to have a degree of arbitrariness in the definition of a PMMs in realistic systems that will always depend on the threshold values chosen in Eq. (The fate of poor man’s Majoranas in the long Kitaev chain limit).
@@
Connection to topology: Although the requirement to define a ROT creates some ambiguity, it is often implicitly assumed that imperfect PMMs in more realistic microscopic models, such as that shown in Fig. 1(a), will still evolve into topological MBSs in the long chain limit, i.e., in the system shown in Fig. 1(b). In the following, we will demonstrate that this is not necessarily the case and show that none of the previously used experimental signatures for PMMs can clearly distinguish between PMMs that would evolve into topological states, from those that evolve into highly localized trivial states.
@@
For instance, in Fig. 2, we consider two sets of parameters, both result in a ROT and therefore correspond to highly localized imperfect PMMs. In the minimal Kitaev chain limit, the ground state energy difference $\Delta E$, the charge difference $\Delta Q_{j}$, and the MP $M_{j}$ (all shown in Fig. 2) give qualitatively similar results for both parameter sets. However, in the long chain limit, we observe that only one of the two parameter sets actually results in MBSs, whereas the other parameter set results in states that are highly localized but trivial, see Fig. 3. We verify this in the infinite chain limit using the topological invariant $Q_{\mathbb{Z}_{2}}$, see SM sm . If $Q_{\mathbb{Z}_{2}}=1$ ($Q_{\mathbb{Z}_{2}}=-1$), the system is trivial (topological). In the trivial case, the highly localized state we observe in the ROT is associated with two doubly degenerate states that split from the bulk states and cross zero energy in the long chain limit. We emphasize that these states become degenerate only in the long chain limit, i.e., in short chains $E_{\mathrm{ex}}>0$. We call these states that appear in a ROT, but that are not connected to topological states in the long chain limit “false PMMs”. We note that we set $U=0$ for the examples shown here, however, false PMMs also exist for finite $U$, see SM sm .
@@
Figure 3: True and false PMMs in the long chain limit. Energy spectrum $E$ (black dots) as a function of the Zeeman energy $\Delta_{Z}$ for the minimal chain [(a) and (e)] and in the long chain limit [(b) and (f)]. The magenta line indicates the topological invariant, calculated in the infinite chain limit, see SM sm . A red (yellow) background in the spectrum indicates that the system is topological (trivial). (c) and (d): Probability density of the two zero-energy states at $\Delta_{Z}/\Delta=0.8$ from (b). The weight on the normal (superconducting) QDs are indicated by black dots (gray crosses). These states are MBSs. (g) and (h): Probability density of two zero-energy states at $\Delta_{Z}/\Delta=0.8$ in (f). Although these states are at zero-energy and mainly localized on the first and last QD, they are trivial and not related to MBSs. Hence, although in the minimal chain [(a) and (e)], both sets of parameters result in qualitatively similar results, only one of the parameter sets is connected to a topological state in the long chain limit. The parameters for (a)-(d) are as for (a)-(c) in Fig. 2 with $\mu/\Delta=0.884$ and $\mu_{M}/\Delta=1.275$. The parameters for (e)-(h) are as for (d)-(f) in Fig. 2 with $\mu/\Delta=0.593$ and $\mu_{M}/\Delta=-3.836$. The long chain consists of 100 normal QDs and 99 QDs with a pairing potential.
@@
Figure 4: Zero-bias conductance signatures of true PMMs versus false PMMs. Normalized zero-energy conductance $G_{LL}$ [(a), (e), (i), (m), (q), and (u)], $G_{LR}$ [(b), (f), (j), (n), (r), and (v)], $G_{RL}$ [panels (c), (g), (k), (o), (s), and (w)], and $G_{RR}$ [(d), (h), (l), (p), (t), and (x)], as $\mu_{L}$ and $\mu_{R}$ are varied independently, for different values of $\mu_{M}$. For each setup all conductance values are normalized with respect to the maximum value. In (a)-(l) the same parameters as for (a)-(c) in Fig. 2 are used, with $\mu_{M}/\Delta=0.975$ for (a)-(d), $\mu_{M}/\Delta=1.275$ (corresponding to the value for the ROT) for (e)-(h), and $\mu_{M}/\Delta=1.975$ for (i)-(l). The coupling to the leads is $t_{l}/\Delta=0.005$. In (m)-(x) the same parameters as for (d)-(f) in Fig. 2 are used, with $\mu_{M}/\Delta=-4.236$ for (m)-(p), $\mu_{M}/\Delta=-3.836$ (corresponding to the value for the ROT) for (q)-(t), and $\mu_{M}/\Delta=-3.336$ for (u)-(x). The coupling amplitude to the leads is $t_{l}/\Delta=0.001$. Although both parameter sets give qualitatively similar results, only the parameters for (a)-(l) are related to true MBSs in the long chain limit, whereas (m)-(x) are related to trivial states.

### v2

@@ Abstract
A minimal Kitaev chain, consisting of two quantum dots connected via a superconductor, can host highly localized near-zero-energy states, known as poor man’s Majoranas (PMMs). These states have been proposed as promising candidates to study Majorana bound states (MBSs) in a highly tunable setup. However, it is unclear whether and how PMMs observed in real systems are actually connected to the topological phase of the full Kitaev chain. Here, we study PMMs using a microscopic model and show that, in the long chain limit, not all PMMs are related to topological states. Rather, in long chains, some PMMs evolve into trivial highly localized low-energy states. We provide an explanation for the occurrence of these states and show that there is no clear conductance signature that is able to distinguish PMMs that evolve into true topological states from PMMs that evolve into trivial states.
@@
Majorana bound states (MBSs) are quasiparticles that emerge at the boundaries or in vortices of topological superconductors Kitaev (2001) and, due to their non-Abelian exchange statistics Ivanov (2001), have been proposed as candidates to store and manipulate quantum information in a fault-tolerant way Kitaev (2003); Nayak et al. (2008); Elliott and Franz (2015). They were prominently described using the Kitaev chain Kitaev (2001), a simple spinless minimal model relying on $p$-wave superconductivity. Nanowires with strong spin-orbit interaction (SOI) proximitized by superconductors have in recent years been prominent systems to realize an effective Kitaev chain Lutchyn et al. (2010); Oreg et al. (2010); Stanescu et al. (2011); Mourik et al. (2012); Das et al. (2012); Deng et al. (2012); Laubscher and Klinovaja (2021). However, disorder has hindered the conclusive observation of MBSs in these heterostructures since disorder can result in signatures that mimic MBSs Kells et al. (2012); Lee et al. (2012); Rainis et al. (2013); Roy et al. (2013); Ptok et al. (2017); Liu et al. (2017); Moore et al. (2018a, b); Reeg et al. (2018); Vuik et al. (2019); Stanescu and Tewari (2019); Woods et al. (2019); Chen et al. (2019); Awoga et al. (2019); Prada et al. (2020); Yu et al. (2021); Das Sarma and Pan (2021); Valentini et al. (2021); Hess et al. (2021, 2023a, 2022). Due to the issues with disorder in nanowires, realizing an effective Kitaev chain as an array of quantum dots (QDs) has been proposed as an alternative platform to host MBSs with a high degree of tunability Sau and Sarma (2012); Leijnse and Flensberg (2012); Fulga et al. (2013); Sothmann et al. (2013); Li et al. (2014); Su et al. (2017); Tsintzis et al. (2022); Liu et al. (2022); Dvir et al. (2023); Sanches et al. (2023); Liu et al. (2023); Koch et al. (2023); Souto et al. (2023); Zatelli et al. (2023); Samuelson et al. (2024); Tsintzis et al. (2024); Pino et al. (2024a); Luna et al. (2024); Bordin et al. (2024a); Boross and Pályi (2024); Liu et al. (2024a); Vilkelis et al. (2024); Ezawa (2024); Bozkurt et al. (2024); van Driel et al. (2024); Pino et al. (2024b); Geier et al. (2024); Ten Haaf et al. (2024); Kocsis et al. (2024); Cayao (2024); Miles et al. (2024); Cayao and Aguado (2024); Liu et al. (2024b); Alvarado et al. (2024); Svensson and Leijnse (2024); Álvaro Gómez-León et al. (2024); Pan et al. (2024); Pandey and Dagotto (2024); Liu et al. (2024c); Benestad et al. (2024); Luethi et al. (2024); ten Haaf et al. (2024). In these proposals, the QDs are separated by superconducting sections, which couple the QDs through elastic cotunneling (ECT) and crossed Andreev reflection (CAR) Recher et al. (2001); Lesovik et al. (2001); Falci et al. (2001); Bouchiat et al. (2002); Feinberg (2003); Hofstetter et al. (2009); Herrmann et al. (2010). In such setups, an inhomogenous magnetic field or SOI can enable effective $p$-wave superconductivity using a conventional $s$-wave superconductor Sau and Sarma (2012); Leijnse and Flensberg (2012).
@@
It has been shown that a fully spin-polarized model consisting of only two QDs with a superconductor transmitting ECT and CAR between the QDs is sufficient to obtain states similar to MBSs Leijnse and Flensberg (2012). These states share most of their properties with MBSs found in long Kitaev chains, i.e., they are at zero energy, separated from excited states by a finite gap, and have non-Abelian exchange statistics. However, they exist only at finely tuned points of parameter space, called “sweet spots”, and therefore they lack topological protection. Due to this lack of protection MBSs in minimal Kitaev chains are often called “poor man’s Majoranas” (PMMs).
@@
Recent demonstrations of control over the strength of ECT and CAR Liu et al. (2022); Wang et al. (2022); Kürtössy et al. (2022); de Jong et al. (2023); Bordin et al. (2023); Wang et al. (2023); Bordin et al. (2024b) have spurred great interest in PMMs. Although originally derived using the spinless Kitaev chain, PMMs have also been studied in spinful models. It was shown in Ref. 74 that sweet spots with perfectly localized zero-energy PMMs can exist in an idealized spinful model, however, reaching these sweet spots requires complete control over many degrees of freedom. In more realistic models and experiments such control is impossible and only “imperfect PMMs” are possible, i.e., states that are close to zero energy and highly – but not perfectly – localized. Nevertheless, the study of imperfect PMMs in minimal Kitaev chains has attracted much attention due to the hope that they are the building blocks of topological states in longer QD chains Sau and Sarma (2012); Leijnse and Flensberg (2012); Dvir et al. (2023); Zatelli et al. (2023); Samuelson et al. (2024); Tsintzis et al. (2024); Pino et al. (2024a); Luna et al. (2024); Bordin et al. (2024a); Ezawa (2024); Miles et al. (2024); Liu et al. (2024b); Alvarado et al. (2024); Svensson and Leijnse (2024). Furthermore, in recent experiments, conductance signatures associated with PMMs have been found Dvir et al. (2023); Ten Haaf et al. (2024); Zatelli et al. (2023); Bordin et al. (2024a); van Driel et al. (2024); ten Haaf et al. (2024). It is therefore important to clarify whether imperfect PMMs evolve into topological or trivial states in the long chain limit.
@@
Figure 1: Microscopic model of a minimal Kitaev chain and the long chain limit. (a): Minimal Kitaev chain, consisting of three QDs. The left and right QDs, representing two Kitaev sites, have a chemical potential $\mu$, Zeeman energy $\Delta_{Z}$, and on-site Coulomb interaction $U$. The middle QD M has a superconducting pairing potential $\Delta$ and, due to the screening by the superconductor, we assume no Zeeman energy or Coulomb interaction. This middle QD models an ABS that transmits the effective couplings between the two outer QDs. The hopping amplitude is $t$ and $U_{\mathrm{SOI}}(\Phi_{\mathrm{SOI}}/2)$ characterizes the SOI. (b): Long chain limit. Normal QDs are white, they are equivalent to the QDs L and R in (a) and each represents one normal site. The QDs with a pairing potential are blue and they are equivalent to the QD M in (a).
@@
In this Letter, we demonstrate that imperfect PMMs in a microscopic model of the minimal Kitaev chain are neither necessary nor sufficient to obtain MBSs in the long chain limit. We explain the origin of these imperfect PMMs and demonstrate that there is no clear conductance signature in a minimal Kitaev chain that can distinguish PMMs that evolve into true MBSs from PMMs that evolve into trivial states.
@@
Model: To model a minimal Kitaev chain microscopically we consider a system of three QDs, see Fig. 1(a), based on the model introduced in Ref. 40. The left and right QDs, modelling the two sites of a minimal Kitaev chain, have chemical potentials $\mu_{L}$ and $\mu_{R}$, respectively (if not stated otherwise, we assume $\mu_{L}=\mu_{R}\equiv\mu$), Zeeman energy $\Delta_{Z}$, and on-site Coulomb repulsion $U$. The middle QD is proximitized by a superconductor and models an Andreev bound state (ABS) that effectively transmits ECT and CAR between the outer QDs Tsintzis et al. (2022); Liu et al. (2022); Miles et al. (2024); Souto et al. (2023); Liu et al. (2024c); Zatelli et al. (2023); Luna et al. (2024); Valentini et al. (2024). The middle QD also causes local Andreev reflection (LAR) and renormalizes other parameters of the outer QDs Luethi et al. (2024). This middle dot has a superconducting pairing potential $\Delta$ and chemical potential $\mu_{M}$. Due to screening by the superconductor we assume that the middle QD has no Zeeman energy and no Coulomb repulsion. We note, however, that it has been shown in Ref. 40 that, even without this assumption, PMMs persist. The hopping between the QDs is determined by the hopping amplitude $t$ and the SOI angle $\Phi_{\mathrm{SOI}}$. The Hamiltonian describing the system is given by
@@
where $n_{j,\sigma}=d_{j,\sigma}^{\dagger}d_{j,\sigma}$, $d_{j,\sigma}^{\dagger}$ ($c_{\sigma}^{\dagger}$) creates a particle on dot $j\in\{L,R\}$ (the middle QD M) and the notation $\sigma\Delta_{Z}$ means $+\Delta_{Z}$ for $\sigma=\uparrow$ and $-\Delta_{Z}$ for $\sigma=\downarrow$. The chemical potentials are measured with respect to the chemical potential of the superconductor. Assuming that SOI is perpendicular to the magnetic field, the SOI matrix is
@@
Figure 2: Quality measures for true PMMs versus false PMMs. Energy difference $\Delta E$ [(a) and (d)], charge difference $\Delta Q_{j}$ [(b) and (e)], and the Majorana polarisation $M_{j}$ [(c) and (f)] for two parameter sets. The area circled in black is the TR. Both sets of parameters lead to qualitatively equal results, and both lead to a TR with highly localized near-zero-energy states, i.e., PMMs. However, only (a)-(c) are connected to true MBSs in the long chain limit, whereas (d)-(f) are connected to trivial states in the long chain limit, see Fig. 3. The parameters for (a)-(c) are $t/\Delta=0.42$, $\Phi_{\mathrm{SOI}}=0.26\pi$, $U=0$, and $\Delta_{Z}/\Delta=0.8$. The largest excitation gap in the TR is $E_{\mathrm{ex}}/\Delta=0.103$. The parameters for (d)-(f) are $t/\Delta=0.99$, $\Phi_{\mathrm{SOI}}=0.44\pi$, $U=0$, and $\Delta_{Z}/\Delta=0.8$. The largest excitation gap in the TR is $E_{\mathrm{ex}}/\Delta=0.084$. The TR threshold values for all panels are $\Delta E_{\mathrm{th}}/\Delta=10^{-3}$, $\Delta Q_{\mathrm{th}}=0.05$, and $M_{\mathrm{th}}=0.05$.
@@
Threshold region: A sweet spot is a point in parameter space that fulfills $\Delta E=0$, $\Delta Q_{L}=\Delta Q_{R}=0$, $M_{L}=M_{R}=1$, and $E_{\mathrm{ex}}>0$. The corresponding “perfect PMMs” to these sweet spots are perfectly localized and at zero energy. It was shown in Ref. Luethi et al. (2024) that, for the microscopic Hamiltonian of a minimal Kitaev chain in Eq. (1), no sweet spots could be found. Instead, the notion of a “threshold region” (TR) was introduced. It is a region of parameter space where the PMM conditions are approximately realized, i.e.,
@@
where $\Delta E_{\mathrm{th}}$, $E_{\mathrm{ex,th}}$, $\Delta Q_{\mathrm{th}}$, and $M_{\mathrm{th}}$ are threshold values which need to be chosen. The corresponding “imperfect PMMs” are highly – but not perfectly – localized near-zero-energy states. This highlights that it is necessary to have a degree of arbitrariness in the definition of PMMs in realistic systems that will always depend on the threshold values chosen in Eq. (The fate of poor man’s Majoranas in the long Kitaev chain limit).
@@
Connection to topology: Although the requirement to define a TR creates some ambiguity, it is often implicitly assumed that imperfect PMMs in more realistic microscopic models, such as that shown in Fig. 1(a), will still evolve into topological MBSs in the long chain limit, i.e., in the system shown in Fig. 1(b). In the following, we demonstrate that this is not necessarily the case and show that none of the previously used experimental signatures for PMMs can clearly distinguish between PMMs that would evolve into topological states, from those that evolve into highly localized trivial states.
@@
For instance, in Fig. 2, we consider two sets of parameters, both result in a TR and therefore corresponding to imperfect PMMs. In the minimal Kitaev chain, the ground state energy difference $\Delta E$, the charge difference $\Delta Q_{j}$, and the MP $M_{j}$ (all shown in Fig. 2) give qualitatively similar results for both parameter sets. Since the topological character of these states cannot be judged in a minimal chain, we extend the setup to a long uniform chain, keeping all parameters unchanged. We observe that the PMMs of Figs. 2(a)-(c) [(d)-(f)] evolve into MBSs (trivial states) and verify this in the infinite chain limit using the topological invariant $Q_{\mathbb{Z}_{2}}$, see SM sm . If $Q_{\mathbb{Z}_{2}}=1$ ($Q_{\mathbb{Z}_{2}}=-1$), the system is trivial (topological). The highly localized trivial state is associated with two doubly degenerate states that split from the bulk states and cross zero energy in the long chain limit. We emphasize that these states become degenerate only in the long chain limit, i.e., in short chains the excitation energy is finite. We call the PMMs that evolve into topological (trivial) states in the long chain limit “true PMMs” (“false PMMs”). False PMMs are unrelated to MBSs in the sense that they do not act as the minimal building block for MBSs, i.e., by extending the minimal chain, the resulting state is trivial. One could tune parameters to bring the long chain into the topological phase, e.g., set $\Delta_{Z}/\Delta=1.05$ in Fig. 3(f). In this case, however, the corresponding states in the minimal chain are no longer PMMs, as their energy is non-zero, see Fig. 3(e). Therefore, a PMM is neither necessary nor sufficient to ensure that the states evolve into MBSs in a long chain. We note that we set $U=0$ here, but false PMMs also exist for finite $U$, see SM sm .
@@
Figure 3: True and false PMMs in the long chain limit. Energy spectrum $E$ (black dots) as a function of the Zeeman energy $\Delta_{Z}$ for the minimal chain [(a) and (e)] and in the long chain limit [(b) and (f)]. The magenta line indicates the topological invariant, calculated in the infinite chain limit, see SM sm . A red (yellow) background in the spectrum indicates that the system is topological (trivial). (c) and (d): Probability density of the two zero-energy states at $\Delta_{Z}/\Delta=0.8$ from (b). The weight on the normal (superconducting) QDs are indicated by black dots (gray crosses). These states are MBSs. (g) and (h): Probability density of two zero-energy states at $\Delta_{Z}/\Delta=0.8$ in (f). Although these states are at zero-energy and mainly localized on the first and last QD, they are trivial and not related to MBSs. Hence, although in the minimal chain [(a) and (e)], both sets of parameters result in qualitatively similar results, only one of the parameter sets is connected to a topological state in the long chain limit. The chain of (f) could be brought into the topological phase by setting $\Delta_{Z}/\Delta=1.05$, but then the states in the minimal chain are no longer PMMs, as their energy is non-zero, see (e). The parameters for (a)-(d) are as for (a)-(c) in Fig. 2 with $\mu/\Delta=0.884$ and $\mu_{M}/\Delta=1.275$. The parameters for (e)-(h) are as for (d)-(f) in Fig. 2 with $\mu/\Delta=0.593$ and $\mu_{M}/\Delta=-3.836$. The long chain consists of 100 normal QDs and 99 QDs with a pairing potential.
@@
Figure 4: Zero-bias conductance signatures of true PMMs versus false PMMs. Normalized zero-energy conductance $G_{LL}$ [(a), (e), (i), (m), (q), and (u)], $G_{LR}$ [(b), (f), (j), (n), (r), and (v)], $G_{RL}$ [panels (c), (g), (k), (o), (s), and (w)], and $G_{RR}$ [(d), (h), (l), (p), (t), and (x)], as $\mu_{L}$ and $\mu_{R}$ are varied independently, for different values of $\mu_{M}$. For each setup all conductance values are normalized with respect to the maximum value. In (a)-(l) the same parameters as for (a)-(c) in Fig. 2 are used, with $\mu_{M}/\Delta=0.975$ for (a)-(d), $\mu_{M}/\Delta=1.275$ (corresponding to the value for the TR) for (e)-(h), and $\mu_{M}/\Delta=1.975$ for (i)-(l). The coupling to the leads is $t_{l}/\Delta=0.005$. In (m)-(x) the same parameters as for (d)-(f) in Fig. 2 are used, with $\mu_{M}/\Delta=-4.236$ for (m)-(p), $\mu_{M}/\Delta=-3.836$ (corresponding to the value for the TR) for (q)-(t), and $\mu_{M}/\Delta=-3.336$ for (u)-(x). The coupling amplitude to the leads is $t_{l}/\Delta=0.001$. Although both parameter sets give qualitatively similar results, only the parameters for (a)-(l) are related to true MBSs in the long chain limit, whereas (m)-(x) are related to trivial states.